# Time depending magnetization of nanoparticles under radiofrequency fields: relaxation time in water for solid-liquid transition

### v1

@@ 5 Discussion and conclusions
We studied the evolution of $\tau$ for a sample of MNPs suspended in water during the solid-liquid transition of the matrix. Several experimental runs were conducted, all yielding consistent results. The obtained cycles are clean and exhibit a monotonic evolution of all their parameters between the solid and liquid states. A detailed study of this evolution will be the focus of future work. The evolution of $\tau$, in particular, shows a clear transition around the melting point of water, from a value of approximately $0.9\times 10^{-7}$ s for temperatures below -5 C to a constant value of approximately $1.4\times 10^{-7}$ s for temperatures above 5 C. This change can be interpreted as a consequence of the alignment of the anisotropy axes of the MNPs with the field $H(t)$: since the sample was frozen without an applied field, the anisotropy axes are randomly oriented in the solid state. When the aqueous matrix becomes liquid, the particles are free to rotate and align their anisotropy axes with the external field, minimizing energy. The size of the MNPs and previously published results indicate that the dominant relaxation mechanism in liquid water is Nèel relaxation de Sousa et al. (2013); Lima et al. (2014), where the magnetic moment flips within the particles following the external field, which in this case oscillates between the two opposite directions of the generating coil axis. The energy barrier for the moment to flip between these two directions is further conditioned by the direction of the particles’ easy magnetization axis with respect to the applied field. When this axis is perpendicular to the field, there is no barrier, while the maximum value is reached when the easy axis is parallel to the field. Thus, the population of MNPs, initially exhibiting a random distribution of easy axis orientations, ends up with all particles aligned in the direction of the field as the matrix becomes liquid, maximizing the effective relaxation time.

### v2

# Time depending magnetization of nanoparticles under radiofrequency fields: experimental relaxation time in water for solid-liquid transition

@@ 5 Discussion and conclusions
We studied the evolution of $\tau$ for a sample of MNPs suspended in water during the solid-liquid transition of the matrix. Several experimental runs were conducted, all yielding consistent results. The obtained cycles are clean and exhibit a monotonic evolution of all their parameters between the solid and liquid states. A detailed study of this process will be the focus of future work. The evolution of $\tau$, in particular, shows a clear transition around the melting point of water, from a value of approximately 90 ns for temperatures below -5 C to a constant value of approximately 140 ns for temperatures above 5 C, much larger than predicted from numerical simulations of similar particles in Davidson et al. (2024). This change can be interpreted as a consequence of the alignment of the anisotropy axes of the MNPs with the field $H(t)$ as proposed in Shah et al. (2015): since the sample was frozen without an applied field, the anisotropy axes are randomly oriented in the solid state. When the aqueous matrix becomes liquid, the particles are free to rotate and align their anisotropy axes with the external field, minimizing energy. The size of the MNPs and previously published results indicate that the dominant relaxation mechanism in liquid water is Nèel relaxation de Sousa et al. (2013); Lima et al. (2014), where the magnetic moment flips within the particles following the external field, which in this case oscillates between the two opposite directions of the generating coil axis. The energy barrier for the moment to flip between these two directions is further conditioned by the direction of the particles’ easy magnetization axis with respect to the applied field. When this axis is perpendicular to the field, there is no barrier, while the maximum value is reached when the easy axis is parallel to the field. Thus, the population of MNPs, initially exhibiting a random distribution of easy axis orientations, ends up with all particles aligned in the direction of the field as the matrix becomes liquid, maximizing the effective relaxation time. These results support the simulations previously published in Chalifour et al. (2021).